4. An aqueous solution of Mg(OH)2 is added to an aqueous solution of HCl to form solid  MgCl2 and H2O. If 40.5g of Mg(OH)2 and 35.0g of HCl are used: 

a) Define limiting reactant 

b) Determine the limiting reactant. 

c) Calculate the mass of MgCl2 produced after the reaction has completed. d) If the percentage yield of MgCl2 is 85%, calculate the actual mass of MgCl2  obtained. 

[ Ar : Mg =24.3 , H=1.0 , O = 16.0 , Cl = 35.5 ] 




1
Expert's answer
2021-10-10T10:24:10-0400

a) The reactant that controls the amount of product able to be produced by a chemical reaction because it is used up completely.

b) Mg(OH)2 + 2HCl → MgCl2 + 2H2O

M(Mg(OH)2) = 58.3 g\mol

M(HCl) = 36.5 g/mol

n(Mg(OH)2) "= \frac{40.5}{58.3} = 0.694 \;mol"

n(HCl) "= \frac{35.0}{36.5} = 0.959 \;mol"

According to the reaction for one mole of Mg(OH)2 we need two moles of HCl. So, HCl is limitting reactant.

c) n(MgCl2) "= \frac{1}{2}n(HCl) = \frac{1}{2} \times 0.959 = 0.479 \;mol"

M(MgCl2) = 95.3 g/mol

m(MgCl2) "= 95.3 \times 0.479 = 45.65 \;g"

d) Proportion:

45.65 g – 100 %

x g – 85 %

"x = \frac{45.65 \times 85}{100} = 38.80 \;g"